# EvoXplain: When Machine Learning Models Agree on Predictions but Disagree on Why Measuring Mechanistic Multiplicity Across Training Runs

### v1

@@ 1 Introduction
Machine learning models are increasingly deployed in domains where explanations matter, including healthcare, public policy, and scientific discovery. Despite the fact that practitioners routinely retrain models during development and performance tuning, interpretability is typically assessed using a single trained instance—or by averaging explanations across runs—rather than by explicitly examining whether explanations themselves remain consistent under retraining. This practice implicitly assumes that high predictive accuracy corresponds to a single, stable way of reasoning (lipton2018myth; rudin2019stop).
There have been sustained challenges to this assumption in the literature. Perhaps the most well-known is the Rashomon effect, which observes that distinct models may achieve similar predictive performance while relying on different internal logics (breiman2001statistical; marx2020predictive). Related work on underspecification highlights that modern learning pipelines often admit many solutions that are indistinguishable in terms of accuracy, leaving optimisation procedures to select arbitrarily among them (d2022underspecification). Other studies have shown that explanations can vary substantially when models are retrained, or when features are correlated or perturbed, including documented failures of attribution methods under retraining, feature overlap, or input transformations (kindermans2017reliability; nogueira2018stability; kumar2020problems; frye2020missingness). Collectively, these results suggest that explanation stability cannot be assumed a priori. However, existing work does not provide a practical way to quantify how much explanatory variability exists in a given setting, nor whether such variability is structured or merely diffuse.
@@
EvoXplain does not aim to decide which explanation is correct. Instead, it provides evidence about whether a model’s explanations are reliable or whether they depend sensitively on the particular optimisation trajectory taken during training. This perspective aligns with a growing view that interpretability should be evaluated empirically rather than assumed, a position increasingly reflected in discussions of responsible AI, model auditing, and governance (arrieta2020explainable; selbst2018intuitive; ecegai2023aiact). By making explanation stability measurable, EvoXplain enables practical questions such as: does this model tell a consistent story, or does its reasoning change across retraining?
@@ 2.1 Model Performance and Interpretability
A large proportion of studies in machine learning assess models primarily through the lens of accuracy. In the real world, interpretability is often categorized based upon the model being used. Linear models and decision trees, for instance, are often tagged as “interpretable” because practitioners are able to read and comprehend the parameters or structures involved. However, some recent contributions to the field argue that readability cannot be equated with the models achieving high predictive accuracy (lipton2018myth). EvoXplain follows the same direction by considering interpretability as something that can be measured and verified rather than assumed.
@@ 2.2 Rashomon Effects and Predictive Multiplicity
Our work relates to the Rashomon effect in that different models can arrive at the same answer and yet rely on totally different paths or reasoning (breiman2001statistical; semengue2019rashomon). Some recent research considers predictive multiplicity under settings of fairness and of unconstrained risk (umbach2023predictive; marx2020predictive; fisher2019all), showing that the same risk scores can be justified through different models (umbach2023predictive). Unlike these approaches, EvoXplain studies multiplicity within the same model class, trained on the same data, to see whether retraining leads to different explanation patterns.
@@ 2.3 Underspecification and Training Variability
Research on underspecification d2022underspecification shows that training pipelines may converge to many possible solutions that perform equally well. EvoXplain complements this work by looking specifically at explanation behaviour. Instead of only asking whether predictions differ, EvoXplain checks whether internal reasoning differs even when predictions stay the same.
@@ 2.4 Feature Importance Instability
Previous studies indicate that SHAP, LIME, and gradients as explanation methods can produce instability due to retraining, dataset perturbations, or correlated features (kindermans2017reliability; nogueira2018stability; kumar2020problems; agarwal2021uncertainty). Most of these studies focus on instability in some specific values. EvoXplain extends local noise analysis by detecting whether instability forms structured groups of explanations and how often each group appears.
2.5 Stability-Aware Interpretability
Certain approaches quantify uncertainty in explanations, such as reporting ranges or using measures of variance (molnar2022interpretable; agarwal2021uncertainty; visani2022statistical). These methods help describe the extent of dispersion across runs, but usually do not assess whether the dispersion pertains to different explanatory modes. EvoXplain works on identifying and quantifying such modes.
@@ 2.6 Mechanistic Interpretability in Deep Models
The understanding of circuits and the internal functioning of large neural networks remains an ongoing area of research (olah2020circuits). Specific techniques tend to rely on handcrafted analyses or other tools engineered for individual designs. On the other hand, EvoXplain focuses on using generic attribution outputs to determine mechanism-level differences.
@@ 3 Problem Formulation
Let ${\theta_{r}}_{r=1}^{R}$ denote $R$ independent training runs on $\mathcal{D}$, and let $\mathbf{e}_{r}=\Phi(\theta_{r})$ denote the corresponding explanation vectors. EvoXplain analyses the empirical distribution of $\mathbf{e}_{r}$} to determine:
@@ 4.7 Algorithm Summary
1:Train $\mathcal{A}$ for $R$ independent seeds; collect attribution vectors $E={\mathbf{e}_{r}}$
2:Normalise: $\hat{\mathbf{e}}_{r}\leftarrow(\mathbf{e}_{r}-\mu)/|\mathbf{e}_{r}-\mu|1$
3:for $k=1$ to $K{\max}$ do
@@
8:Assign each run to cluster $c_{r}\in{1,\dots,k^{\star}}$
@@
11:return $(k^{\star},H,{p_{j}},{\bar{\mathbf{e}}_{j}})$
@@ 4.11 Diagnostic Applications
EvoXplain can serve as a screening tool for interpretability audits, fairness assessments, scientific analysis, and model governance. By identifying cases where explanations fail to stabilise, it signals when explanation use or aggregation should be treated with caution. In high-stakes domains such as healthcare, criminal risk assessment, and public policy, EvoXplain can highlight when additional scrutiny or domain-specific review is warranted.
@@ 7.1.1 Breast Cancer
We start with Logistic Regression on the Breast Cancer dataset, since this is the simplest model in our study and, in principle, one where we would expect explanations to be stable. When the regularisation strength is allowed to vary, this expectation does not hold.
@@
For a representative split (split seed 104), we analyse 1,000 independent training runs. In explanation space, the SHAP vectors do not form a single cloud. Instead, they organise along a smooth, low-dimensional manifold with clear separation between regions (Fig. 1). Clustering in explanation space yields an optimal solution at k=2, with a silhouette score of 0.608 (compared to 0.566 for k=3 and steadily decreasing thereafter), indicating well-separated explanation modes. The normalized explanation entropy is 0.991 (maximum possible value 1.0), showing that runs are almost evenly distributed across the two modes rather than collapsing into one dominant explanation.
[…]
@@
These differences are reflected in the explanations themselves. In split 104, the dominant features in Cluster 0 are classical size-related measurements such as radius error, worst radius, and worst area, whereas Cluster 1 places substantially more weight on error-based and shape-related features such as fractal dimension error and area error, with absolute SHAP magnitudes several times larger than those seen in Cluster 0. Thus, the model is not merely reweighing the same features slightly differently, but relying on qualitatively different explanatory patterns.
@@
Importantly, the average attribution vector—computed as the arithmetic mean of all SHAP vectors across runs—does not lie on the populated manifold. In the split-level embedding, it falls in a low-density region between clusters (Fig. 1), indicating that this average explanation does not correspond to a typical trained model instance.
@@
When we stack explanation vectors across all splits, the same qualitative picture persists. The universal explanation manifold shows multiple distinct tracks that remain separated across resampled datasets, while predictive accuracy remains high along all tracks (Fig. 2). The global average attribution vector, computed across all runs and all clusters, again lies away from the high-density regions of the manifold, reinforcing that averaging explanations across stochastic runs produces a vector that is not representative of deployed models.
@@
To assess whether these explanation differences can translate into different predictions at the individual level, we compare representative models from the two clusters in split 104. Across held-out test instances, the mean absolute disagreement in predicted malignant probability is 0.049, but the tail of the distribution is substantial, with a maximum disagreement of 0.848. For example, for test instance 30 (ground truth: benign), one representative model assigns a malignant probability of 0.15, while the other assigns 0.999. Similar large disagreements are observed for multiple benign and malignant cases, with disagreement scores exceeding 0.5 for several individuals. These discrepancies arise despite both models achieving high overall accuracy.
Control Experiment: Stability under Fixed Regularization.
@@
As a control, we repeat the same analysis under fixed regularisation strength ($C=1.0$). Across ten independent train-test splits, each with 1000 runs, the explanation vectors consistently collapse to a single cluster ($k=1$), with negligible dispersion in explanation space. This confirms that the observed explanation multiplicity is not an artefact of the SHAP method or dimensionality reduction, but emerges specifically when the training process admits multiple equally accurate models.
@@
Figure 1: Breast Cancer — Logistic Regression. Left: PCA embedding of explanation (attribution) vectors obtained from repeated training runs on the same train–test split. Points are coloured by cluster assignment identified in explanation space. Red stars indicate representative cluster centroids, while the black cross denotes the average attribution vector across all runs. Right: The same embedding with predictive accuracy overlaid. Accuracy remains high across the full extent of the explanation manifold, including across distinct clusters, indicating that multiple explanatory modes coexist at comparable performance.
Figure 2: Breast Cancer - Logistic Regression: universal explanation manifold. PCA embedding of explanation (SHAP) vectors obtained by stacking runs across all train-test splits. Each point corresponds to a trained model instance and is coloured by predictive accuracy. Multiple persistent explanation tracks are visible across splits, indicating that explanation variability is not confined to a single resampling of the data. The black $\times$ marks the global average attribution vector, computed as the arithmetic mean of all SHAP vectors across runs and clusters; this point does not correspond to a typical location on the populated manifold.
@@
Table 1: Split-level summary for Logistic Regression on Breast Cancer under stochastic training. For each train-test split, we report the number of runs, the selected number of explanation clusters $k^{*}$, the silhouette score at $k^{*}$, the normalized cluster entropy $H_{\mathrm{norm}}$, and predictive accuracy (mean $\pm$ std across runs). This table quantifies split-to-split consistency of the clustering outcome and confirms that multiplicity is not driven by unstable performance.
Split Runs $k^{*}$ Silh. $(k^{*})$ $H_{\mathrm{norm}}$ Acc. mean Acc. std
100 1000 3 0.554 0.972 0.973 0.004
101 1000 2 0.603 1.000 0.960 0.008
102 1000 3 0.540 0.982 0.971 0.010
103 1000 2 0.585 1.000 0.977 0.012
104 1000 2 0.608 0.991 0.967 0.010
105 1000 2 0.604 0.985 0.956 0.005
106 1000 2 0.647 0.997 0.967 0.009
107 1000 3 0.562 0.983 0.958 0.005
108 1000 2 0.616 0.992 0.969 0.009
109 1000 2 0.603 0.984 0.963 0.007
@@
Table 2: Cluster-level summary for Logistic Regression on Breast Cancer under stochastic training. Each row corresponds to one explanation cluster within one train–test split. We report cluster size, regularisation strength statistics (mean $\pm$ std and range of $C$ values across runs assigned to the cluster), and predictive accuracy (mean $\pm$ std). This table characterises how explanatory basins align with distinct regularisation regimes while maintaining similar predictive performance.
@@
100 0 293 3.275 2.418 [0.639, 9.4] 0.974 0.003
100 1 451 0.151 0.161 [0.010, 0.6] 0.974 0.005
100 2 256 38.547 25.163 [9.528, 100.0] 0.971 0.002
101 0 504 0.221 0.259 [0.010, 1.0] 0.964 0.009
101 1 496 21.742 25.113 [1.042, 100.0] 0.957 0.004
102 0 275 36.484 25.431 [7.997, 100.0] 0.963 0.004
102 1 429 0.129 0.132 [0.010, 0.5] 0.972 0.011
102 2 296 2.726 2.031 [0.522, 7.9] 0.977 0.007
103 0 498 0.212 0.245 [0.010, 1.0] 0.976 0.015
103 1 502 21.494 25.064 [0.986, 100.0] 0.979 0.006
104 0 556 0.325 0.410 [0.010, 1.7] 0.971 0.012
104 1 444 24.132 25.496 [1.684, 100.0] 0.963 0.006
105 0 573 0.369 0.477 [0.010, 2.0] 0.956 0.005
105 1 427 25.021 25.599 [1.969, 100.0] 0.955 0.004
106 0 532 0.272 0.332 [0.010, 1.3] 0.971 0.007
106 1 468 22.972 25.330 [1.349, 100.0] 0.962 0.008
107 0 283 35.670 25.519 [7.428, 100.0] 0.953 0.000
107 1 427 0.127 0.130 [0.010, 0.5] 0.961 0.005
107 2 290 2.575 1.883 [0.513, 7.4] 0.957 0.003
108 0 449 23.882 25.463 [1.608, 100.0] 0.968 0.005
108 1 551 0.313 0.392 [0.010, 1.6] 0.969 0.011
109 0 574 0.372 0.481 [0.010, 2.0] 0.965 0.005
109 1 426 25.075 25.604 [1.987, 100.0] 0.959 0.008
[…]
At the same time, predictive performance remains stable. Mean accuracy across splits ranges from 0.956 to 0.977, with standard deviations typically below 0.012. There is no split in which increased explanation multiplicity coincides with degraded or unstable accuracy, consistent with the split-level manifolds where distinct explanation modes occupy similar performance bands.
[…]
Clusters are not small or degenerate. Cluster sizes are substantial and often roughly balanced (for example, split 104 yields clusters of size 556 and 444). Predictive accuracy remains similar across clusters within each split, with differences typically on the order of a few percentage points or less and overlapping accuracy distributions throughout.
Therefore, Tables 2 and 1 show that explanation multiplicity in Logistic Regression on Breast Cancer is consistent across splits, involves large fractions of runs, and aligns with distinct regularisation regimes, while predictive performance remains uniformly high. As a control, when regularisation strength is fixed, explanations collapse to a single cluster (k=1) across all splits and runs (See code repository), confirming that the observed multiplicity arises specifically when the training process admits multiple equally accurate solutions.

### v2

@@ 1 Introduction
Machine learning models are increasingly deployed in domains where explanations matter, including healthcare, public policy, and scientific discovery. Despite the fact that practitioners routinely retrain models during development and performance tuning, interpretability is typically assessed using a single trained instance—or by averaging explanations across runs—rather than by explicitly examining whether explanations themselves remain consistent under retraining. This practice implicitly assumes that high predictive accuracy corresponds to a single, stable way of reasoning (Lipton (2018); Rudin (2019)).
There have been sustained challenges to this assumption in the literature. Perhaps the most well-known is the Rashomon effect, which observes that distinct models may achieve similar predictive performance while relying on different internal logics (Breiman (2001); Marx et al. (2020)). Related work on underspecification highlights that modern learning pipelines often admit many solutions that are indistinguishable in terms of accuracy, leaving optimisation procedures to select arbitrarily among them (D’Amour et al. (2022)). Other studies have shown that explanations can vary substantially when models are retrained, or when features are correlated or perturbed, including documented failures of attribution methods under retraining, feature overlap, or input transformations (Kindermans et al. (2017); Nogueira et al. (2018); Kumar et al. (2020); Frye et al. (2020)). Collectively, these results suggest that explanation stability cannot be assumed a priori. However, existing work does not provide a practical way to quantify how much explanatory variability exists in a given setting, nor whether such variability is structured or merely diffuse.
@@
EvoXplain does not aim to decide which explanation is correct. Instead, it provides evidence about whether a model’s explanations are reliable or whether they depend sensitively on the particular optimisation trajectory taken during training. This perspective aligns with a growing view that interpretability should be evaluated empirically rather than assumed, a position increasingly reflected in discussions of responsible AI, model auditing, and governance (Arrieta and others (2020); Selbst and Barocas (2018); European Commission (2023)). By making explanation stability measurable, EvoXplain enables practical questions such as: does this model tell a consistent story, or does its reasoning change across retraining?
@@ 2.1 Model Performance and Interpretability
A large proportion of studies in machine learning assess models primarily through the lens of accuracy. In the real world, interpretability is often categorized based upon the model being used. Linear models and decision trees, for instance, are often tagged as “interpretable” because practitioners are able to read and comprehend the parameters or structures involved. However, some recent contributions to the field argue that readability cannot be equated with the models achieving high predictive accuracy (Lipton (2018)). EvoXplain follows the same direction by considering interpretability as something that can be measured and verified rather than assumed.
@@ 2.2 Rashomon Effects and Predictive Multiplicity
Our work relates to the Rashomon effect in that different models can arrive at the same answer and yet rely on totally different paths or reasoning (Breiman (2001); Semengue et al. (2019)). Some recent research considers predictive multiplicity under settings of fairness and of unconstrained risk (Umbach et al. (2023); Marx et al. (2020); Fisher et al. (2019)), showing that the same risk scores can be justified through different models (Umbach et al. (2023)). Unlike these approaches, EvoXplain studies multiplicity within the same model class, trained on the same data, to see whether retraining leads to different explanation patterns.
@@ 2.3 Underspecification and Training Variability
Research on underspecification D’Amour et al. (2022) shows that training pipelines may converge to many possible solutions that perform equally well. EvoXplain complements this work by looking specifically at explanation behaviour. Instead of only asking whether predictions differ, EvoXplain checks whether internal reasoning differs even when predictions stay the same.
@@ 2.4 Feature Importance Instability
Previous studies indicate that SHAP, LIME, and gradients as explanation methods can produce instability due to retraining, dataset perturbations, or correlated features (Kindermans et al. (2017); Nogueira et al. (2018); Kumar et al. (2020); Agarwal et al. (2021)). Most of these studies focus on instability in some specific values. EvoXplain extends local noise analysis by detecting whether instability forms structured groups of explanations and how often each group appears.
2.5 Rashomon-Set Explanations in AutoML
A related line of work incorporates model multiplicity into post-hoc explanations by operating over Rashomon sets—collections of near-optimal models within a tolerance of the best model’s loss. Cavus et al. (2025) propose the Rashomon Partial Dependence Profile (Rashomon PDP), which aggregates partial dependence profiles across the Rashomon set and reports uncertainty bands to highlight regions of disagreement. While such approaches usefully acknowledge that multiple near-optimal models exist, they typically treat multiplicity as uncertainty around a single aggregated explanation. EvoXplain studies a different phenomenon: structured, discrete explanatory modes (basins) among equally accurate models, and shows that explanation aggregation can yield a ghost explanation that may not correspond to any realised model instance. In other words, rather than summarising variability around one explanation, EvoXplain detects and quantifies distinct mechanistic regimes within the same training pipeline.
@@ 2.6 Mechanistic Interpretability in Deep Models
The understanding of circuits and the internal functioning of large neural networks remains an ongoing area of research (Olah et al. (2020)). Specific techniques tend to rely on handcrafted analyses or other tools engineered for individual designs. On the other hand, EvoXplain focuses on using generic attribution outputs to determine mechanism-level differences.
2.7 Explainable AI for Science and Metafeature-Space Diagnostics
Recent work in explainable AI for scientific workflows studies where a fixed model tends to improve or degrade, by searching for subgroups of the feature space associated with changes in predictive behaviour. Soares et al. introduce Meta Subspace Analysis (MSA), which discovers exceptional subspaces and relates their metafeatures (e.g., distributional and dependency summaries) to model (mis)behavior via association rules (Soares et al. (2025)). This perspective is complementary to EvoXplain. MSA characterises heterogeneity across data subspaces for a given trained model, whereas EvoXplain characterises heterogeneity across trained model instances under the same pipeline. Thus, MSA does not overlap with EvoXplain’s core object of study (explanation basins over retraining), but motivates future combinations, such as measuring explanation multiplicity within subspaces identified as behaviourally exceptional.
2.8 Explanation Stability in Safety-Critical Machine Learning
Beyond interpretability research, the safety assurance community has emphasised the need for explicit evidence about model robustness under retraining. Habli and colleagues propose structured assurance methodologies for machine learning components, such as AMLAS and the “BIG Argument” for AI safety (Habli et al. (2025); Hawkins et al. (2021)), which require that system behaviour be justified and verified beyond predictive accuracy alone. EvoXplain contributes to this assurance perspective by offering a concrete diagnostic for explanation stability. When explanations diverge across retrainings of equally accurate models, this reflects a form of epistemic uncertainty that must be accounted for in any rigorous ML safety case. In this sense, EvoXplain does not replace assurance frameworks, but provides a measurable source of evidence that can be incorporated within them.
@@ 3 Problem Formulation
Let $\{\theta_{r}\}_{r=1}^{R}$ denote $R$ independent training runs on $\mathcal{D}$, and let $\mathbf{e}_{r}=\Phi(\theta_{r})$ denote the corresponding explanation vectors. EvoXplain analyses the empirical distribution of $\{\mathbf{e}_{r}\}$ to determine:
@@ 4.7 Algorithm Summary
1:Train $\mathcal{A}$ for $R$ independent seeds; collect attribution vectors $E=\{\mathbf{e}_{r}\}$
2:Normalise: $\hat{\mathbf{e}}_{r}=\frac{\mathbf{e}_{r}-\mu}{\|\mathbf{e}_{r}-\mu\|_{2}}$
3:for $k=1$ to $K_{\max}$ do
@@
8:Assign each run to cluster $c_{r}\in\{1,\dots,k^{\star}\}$
@@
11:return $(k^{\star},H,\{p_{j}\},\{\bar{\mathbf{e}}_{j}\})$
@@ 4.11 Diagnostic Applications
EvoXplain can serve as a screening tool for interpretability audits, fairness assessments, scientific analysis, and model governance. By identifying cases where explanations fail to stabilise, it reveals when explanation use or aggregation should be treated with caution. In critical domains such as healthcare, criminal risk assessment, and public policy, EvoXplain can highlight when additional scrutiny or domain-specific review is warranted.
@@ 7.1.1 Breast Cancer
We begin with Logistic Regression on the Breast Cancer dataset, the simplest model in our study and, in principle, one where explanations should be stable. When regularisation strength is allowed to vary, this expectation does not hold.
@@
For a representative split (split seed 109), we analyse 1,000 independent training runs. In explanation space, the SHAP vectors do not form a single cloud. Instead, they organise along a smooth, low-dimensional arc with clear separation between regions (Fig. 1). Clustering in explanation space yields an optimal solution at k=2k{=}2, with a silhouette score of 0.776, indicating well-separated explanation modes. The normalised explanation entropy is 0.922, showing that runs are distributed across the two modes rather than collapsing into a single dominant explanation.
The two clusters are not small perturbations of one another. Cluster 0 contains 663 runs with mean $C\approx 10.84$, while Cluster 1 contains 337 runs with mean $C\approx 12.20$. Despite occupying different regions of the explanation manifold, both clusters achieve comparable predictive accuracy (approximately 0.96). The accuracy distributions overlap substantially (Fig. 1, right panel). In other words, both explanation modes achieve similarly high predictive performance.
The clusters correspond to distinct regularisation regimes. The right panel of Fig. 1 shows a smooth gradient of $C$ values along the arc: strong regularisation (low $C$) concentrates at one end, while weak regularisation (high $C$) occupies the other. This gradient structure indicates that explanation geometry is tightly coupled to the regularisation–Loss landscape interaction.
@@
These differences are reflected in the explanations themselves. The dominant features differ qualitatively between clusters. One mode emphasises classical size-related measurements, while the other places substantially more weight on variability- and shape-related features. The model is not merely reweighting the same features slightly differently; it is relying on qualitatively different explanatory patterns.
@@
Importantly, the average attribution vector—computed as the arithmetic mean of all SHAP vectors across runs—does not lie on the populated manifold. In the split-level embedding, it falls in a sparsely populated region between clusters (Fig. 1), indicating that this average explanation does not correspond to a typical trained model instance.
@@
When explanation vectors are stacked across all ten splits, the same qualitative picture persists. The universal explanation manifold shows ten parallel arcs, each corresponding to a different train–test realisation, that remain separated in explanation space, while predictive accuracy remains uniformly high along all tracks (Fig. 2). The global average attribution vector, computed across all runs and all splits, lies again away from the high-density regions of the manifold. This reinforces that averaging explanations across stochastic runs produces a vector that is not representative of any deployed model instance.
@@
To assess whether these explanation differences translate into different predictions at the individual level, we compare representative models from the two clusters in split 109. In all held test instances, the mean absolute disagreement in the predicted malignant probability is 0.067 and the median is 0.035. However, the tail of the distribution is substantial: the maximum disagreement reaches $\Delta=0.73$, with 34 instances exceeding $\Delta>0.1$, 12 instances exceeding $\Delta>0.2$, and 4 instances exceeding $\Delta>0.3$. For the highest-disagreement case, one representative model assigns a malignant probability substantially different from the other, despite both models achieving high overall accuracy. These discrepancies arise even though aggregate performance metrics are indistinguishable.
@@
Control Experiment: Stability under Fixed Regularisation. As a control, we repeat the same analysis under fixed regularisation strength ($C=1.0$). Across ten independent train–test splits, each with 1,000 runs, the explanation vectors consistently collapse to a single cluster ($k=1$), with negligible dispersion in explanation space. This confirms that the observed explanation multiplicity is not an artefact of the SHAP method or dimensionality reduction, but emerges specifically when the training process admits multiple equally accurate models.
@@
Figure 1: Breast Cancer - Logistic Regression (split 109). Left: PCA embedding of explanation (SHAP) vectors obtained from 1,000 independent training runs on a fixed train-test split. Points are coloured by cluster assignment identified in normalised explanation space. Silhouette-based selection yields $k=2$ clusters with normalised entropy $H_{\mathrm{norm}}=0.922$, indicating near-uniform distribution across explanatory modes. Cluster 0 (blue, $n=663$, mean $C=10.84$) and Cluster 1 (orange, $n=337$, mean $C=12.20$) occupy distinct regions of the explanation manifold. Stars indicate cluster centroids; the black cross denotes the average attribution vector across all runs, which lies in a sparsely populated region between clusters. Right: The same embedding with regularisation strength $C$ overlaid. The smooth colour gradient along the arc reveals that explanation geometry is strongly coupled to the regularisation regime, with low-$C$ (strong regularisation) and high-$C$ (weak regularisation) runs occupying opposite ends of the manifold. Despite this mechanistic separation, both clusters achieve comparable predictive accuracy ($\approx 0.96$).
Figure 2: Breast Cancer - Logistic Regression: universal explanation manifold. PCA embedding of SHAP vectors obtained by stacking N=10, [^0] N=10{,}000 training runs across 10 independent train–test splits. Top-left: Points coloured by split seed reveal 10 parallel arcs in explanation space, each corresponding to a different data realisation. The consistent arc geometry across splits indicates that explanation multiplicity is a stable property of the model class rather than an artefact of a particular train–test partition. Top-right: Accuracy overlay shows uniformly high predictive performance (0.94–0.99) across the entire manifold, confirming that distinct explanatory modes coexist at comparable accuracy. Bottom-left: Raw regularisation strength $C$ varies smoothly along each arc. Bottom-right: $\log_{10}(C)$ overlay with diverging colourmap clearly delineates the two regularisation regimes: strong regularisation (blue, $\log_{10}C<0$) concentrates at the left endpoint of each arc, while weak regularisation (red, $\log_{10}C>0$) occupies the right endpoint. The first principal component captures 69.4% of variance and aligns with the regularisation gradient, while PC2 (7.0%) separates individual splits. This visualisation demonstrates that mechanistic multiplicity persists across data resampling and is governed by the regularisation–Loss landscape interaction rather than by noise or overfitting.
@@
Table 1: Split-level summary for Logistic Regression on Breast Cancer under stochastic training with within-split clustering. For each train–test split, we report the number of runs, the selected number of explanation clusters $k$, the silhouette score at $k$, the normalised cluster entropy $H_{\mathrm{norm}}$, and predictive accuracy (mean $\pm$ std across runs). All splits consistently yield $k=2$ explanation clusters with high silhouette scores (0.76–0.82) and near-maximal entropy (0.91–0.94), indicating stable, well-separated, and balanced mechanistic multiplicity across data realisations.
Split Runs $k$ Silh. ($k$) $H_{\mathrm{norm}}$ Acc. mean Acc. std
100 1000 2 0.771 0.919 0.974 0.004
101 1000 2 0.795 0.943 0.961 0.007
102 1000 2 0.757 0.925 0.971 0.010
103 1000 2 0.774 0.932 0.976 0.012
104 1000 2 0.770 0.929 0.968 0.010
105 1000 2 0.806 0.944 0.956 0.005
106 1000 2 0.817 0.912 0.967 0.009
107 1000 2 0.761 0.920 0.957 0.005
108 1000 2 0.791 0.938 0.969 0.008
109 1000 2 0.776 0.922 0.962 0.007
@@
Table 2: Cluster-level summary for Logistic Regression on Breast Cancer under stochastic training with within-split clustering. Each row corresponds to one explanation cluster within one train–test split. We report cluster size, regularisation strength statistics (mean $\pm$ std and range of $C$ values across runs assigned to the cluster), and predictive accuracy (mean $\pm$ std). Notably, both clusters within each split span the full range of $C$ values with similar means, indicating that the mechanistic separation in explanation space is not simply a partition by regularisation strength. The clusters represent genuinely distinct explanatory modes that emerge across the entire hyperparameter range.
@@
100 0 666 10.139 19.776 [0.010, 99.4] 0.974 0.004
100 1 334 10.110 19.318 [0.010, 96.3] 0.974 0.004
101 0 640 10.011 19.784 [0.010, 99.9] 0.961 0.007
101 1 360 14.598 23.484 [0.010, 95.0] 0.960 0.007
102 0 660 11.190 20.888 [0.010, 99.6] 0.971 0.010
102 1 340 11.889 20.636 [0.010, 99.1] 0.971 0.009
103 0 652 10.487 20.645 [0.010, 99.3] 0.976 0.012
103 1 348 12.280 21.910 [0.011, 99.0] 0.976 0.011
104 0 656 11.389 21.649 [0.010, 99.2] 0.969 0.010
104 1 344 11.722 20.984 [0.010, 99.8] 0.968 0.010
105 0 638 10.422 20.393 [0.010, 97.1] 0.956 0.005
105 1 362 11.910 20.600 [0.010, 92.1] 0.955 0.005
106 0 673 10.826 20.983 [0.010, 98.9] 0.967 0.009
106 1 327 10.000 19.920 [0.010, 97.8] 0.968 0.009
107 0 665 10.796 19.679 [0.010, 97.2] 0.958 0.005
107 1 335 11.651 21.597 [0.011, 94.0] 0.957 0.005
108 0 645 10.157 20.061 [0.010, 98.6] 0.970 0.008
108 1 355 10.767 20.794 [0.010, 99.7] 0.969 0.008
109 0 663 10.844 20.565 [0.010, 99.7] 0.962 0.007
109 1 337 12.199 21.267 [0.011, 98.3] 0.962 0.007
Across splits, Logistic Regression on the Breast Cancer dataset consistently exhibits two well-separated explanation basins under stochastic training (Table 1). In every split, clustering selects $k=2$ explanation modes with high silhouette scores (0.76–0.82) and near-maximal normalised entropy (0.91–0.94), indicating that runs are distributed across multiple explanatory mechanisms rather than collapsing into a single dominant solution. This structure is stable across independent train–test realisations.
Importantly, predictive performance remains uniformly high across splits and clusters. Mean accuracy ranges from 0.956 to 0.976, with small standard deviations, and accuracy distributions overlap substantially between explanation basins within each split. There is no split in which explanation multiplicity coincides with degraded or unstable predictive performance, consistent with the split-level and universal explanation manifolds.
Cluster-level analysis (Table 2) further shows that the two explanation basins are not reducible to a simple partition by regularisation strength. Within each split, both clusters span the full sampled range of $C$ values, with similar mean and variance. This indicates that the mechanistic separation observed in explanation space reflects genuinely distinct explanatory modes rather than trivial hyperparameter binning. Note that while position along the arc correlates smoothly with regularisation strength (Fig.1, right panel), cluster membership does not partition by C: runs with identical regularisation can land in different basins depending on other aspects of the optimisation trajectory. As a control, when regularisation strength is fixed, explanations collapse to a single cluster across all splits, confirming that explanation multiplicity emerges specifically when the training process admits multiple equally accurate solutions.